# Autonomous Navigation and Station-Keeping on Near-Rectilinear Halo Orbits

### v1

@@ Abstract
This article develops an optical navigation (OPNAV) and station-keeping pipeline for the near-rectilinear halo orbit (NRHO) in high-fidelity ephemeris model dynamics. The pipeline involves synthetic images used by the non-iterative horizon-based OPNAV algorithm, fed into an extended Kalman filter. The state estimate is used by a controller to maintain the spacecraft’s motion within the vicinity of a reference NRHO. We study differential correction-based and minimization-based implementations of the $x$-axis crossing control scheme, and propose an improved targeting prediction scheme by incorporating the filter’s state covariance with an unscented transform. We also introduce a hysteresis mechanism, which improves station-keeping cost and provides insight into the difference in performance between the differential correction-based and minimization-based approaches. We perform Monte-Carlo experiments to assess the pipeline’s tracking and $\Delta V$ performances. We report several key findings, including the variability of the filter performance with the sensor field of view and measurement locations, station-keeping cost reduction achieved by the unscented transform-based prediction and hysteresis, as well as variability of the cumulative $\Delta V$ as a function of maneuver location due to the periodic structure in the OPNAV-based filter’s estimation accuracy.
@@
$\mathrm{FOV}$  = field of view,  $\text{\,}\mathrm{d}\mathrm{e}\mathrm{g}$
@@
$\boldsymbol{y}$  = measurement
@@
$\boldsymbol{\theta}$  = spacecraft inertial state vector
@@ 1 Introduction
A key challenge of operating space assets in cislunar space is the vast distance from the Earth. The distance makes frequent communication and navigational updates to cislunar spacecraft a challenge due to the scarcity and congestion of ground-based infrastructures, necessitating autonomous capability for guidance, navigation, and control (GNC). A particular orbit of interest is the 9:2 resonant near-rectilinear halo orbit (NRHO) about the Earth-Moon L2 point [1, 2, 3], due to its planned use for the upcoming lunar space station, or Gateway, humanity’s next outpost in space.
@@
Regular station-keeping (SK) maneuvers, also referred to as orbit maintenance maneuvers, are required for spacecraft flying along the NRHO due to the orbit’s unstable modes. SK is required at an appropriate cadence in order to cancel the unstable modes in the presence of errors in state estimation, dynamics model, and control. As traffic on the NRHO is expected to increase, autonomous GNC along this orbit is of critical importance, both to reduce reliance on ground-based navigation updates in nominal operations contexts, as well as to ensure a reliable onboard GNC capability in case of off-nominal scenarios such as communication failures. Today, with the lack of a dedicated cislunar position, navigation, and timing (PNT) architecture, horizon-based optical navigation (OPNAV) [4, 5, 6] is a suitable measurement source that may be acquired and processed onboard a spacecraft along the NRHO. A number of prior simulation-based works considered autonomous navigation using horizon-based OPNAV on lunar libration point orbits (LPOs) [7, 8], as well as specifically on the NRHO [9, 10], but without consideration of SK activities. Several recent missions have successfully demonstrated horizon-based OPNAV for orbit determination in cislunar flight: the Orion spacecraft during the Artemis I mission [11], the LONEStar experiment during the Lunar Flashlight mission [12], and the CAPSTONE mission [13].
@@
There has been significant focus on SK algorithms and their performance under orbit determination error; Shirobokov et al. [14] provides a survey of the various SK schemes and for LPOs in various dynamical regimes. Oguri [15] demonstrates navigation with numerically simulated horizon-based OPNAV and SK based on an optimization-based pre-computed feedback policy. In the context of SK for NRHOs, the $x$-axis crossing control is known to perform favorably [16, 17, 18, 19]; in the recent CAPSTONE mission, this strategy was successfully deployed on the NRHO [20], and has been chosen as the approach for the Gateway [19].
In this work, we evaluate the feasibility and performance of autonomous GNC along NRHO by developing an end-to-end pipeline that consists of a navigation filter relying solely on measurements from processing realistic synthetic images of the Moon, and a SK algorithm that use the recursively filtered state estimate to compute control actions. The synthetic images are generated using a rendering setup previously developed by the authors [21], and are processed to generate position measurements using the state-of-the-art non-iterative horizon-based OPNAV algorithm [4, 5]. We augment the fidelity of the measurement covariance expression of the horizon-based OPNAV algorithm by accounting for the attitude error of the camera. We compare two variants of the $x$-axis crossing control, namely the traditional differential correction-based approach [17, 16, 19] and the optimization-based approach [22]. In addition, we propose an unscented transform (UT)-based targeting scheme that uses the filter’s state covariance information to improve the state prediction. We also introduce hysteresis by setting a higher state violation threshold for executing control than the state violation threshold that the $x$-axis crossing control aims to satisfy. Both the UT-based targeting and hysteresis can be readily implemented within the $x$-axis crossing control scheme, and are found to result in tangible cumulative SK $\Delta V$ cost reduction.
The GNC pipeline is used to conduct multiple Monte-Carlo experiments in which measurement collection and controller hyperparameters are varied. We report the levels of navigation accuracy that can be achieved by horizon-based OPNAV, and quantify its limitations. We compare differential correction-based and minimization-based implementations of the $x$-axis crossing control and show the significant performance difference that arises due to the myopic per-maneuver SK optimization as an approximation of the infinite-horizon cumulative SK cost. Finally, using the pipeline, we study the sensitivity of the cumulative SK cost to SK maneuver location along the NRHO.
@@ 2.1 Equations of Motion
$\dot{\boldsymbol{\theta}}=\begin{bmatrix}\dot{\boldsymbol{r}}\\ \dot{\boldsymbol{v}}\end{bmatrix}=\boldsymbol{f}(t,\boldsymbol{\theta})=\begin{bmatrix}\boldsymbol{v}\\ -\dfrac{\mu}{r^{3}}\boldsymbol{r}+\boldsymbol{a}_{\mathrm{J2}}(t,\boldsymbol{r})+\sum_{i}\boldsymbol{a}_{N_{i}}(t,\boldsymbol{r})+\boldsymbol{a}_{\mathrm{SRP}}(t,\boldsymbol{r})\end{bmatrix}$, (1)
where $\boldsymbol{r}\in\mathbb{R}^{3}$ is the position vector of the spacecraft with respect to an unforced particle at the origin of $\mathcal{F}_{\rm I}$, $\boldsymbol{v}\triangleq\dot{\boldsymbol{r}}\in\mathbb{R}^{3}$ is the rate of change of $\boldsymbol{r}$ in $\mathcal{F}_{\rm I}$, and $\boldsymbol{\theta}^{T}=[\boldsymbol{r}^{T},\boldsymbol{v}^{T}]\in\mathbb{R}^{6}$ is the state vector. In the acceleration terms on the right-hand side, $r=\|\boldsymbol{r}\|$ is the distance from the center of the Moon to the spacecraft, $\mu$ is the standard gravitational parameter of the Moon, $\boldsymbol{a}_{\mathrm{J2}}$ is the perturbing acceleration due to the J2 coefficient of the lunar spherical harmonic gravitational model, $\boldsymbol{a}_{N_{i}}$ is the third-body perturbing acceleration of body $i$, and $\boldsymbol{a}_{\mathrm{SRP}}$ is the perturbing acceleration due to SRP. In this work, we consider third-body perturbations due to the Earth and the Sun. The perturbation terms are given by
@@
$\boldsymbol{a}_{N_{i}}(t,\boldsymbol{r})$ $=-\mu_{i}\left(\dfrac{\boldsymbol{r}-\boldsymbol{d}_{i}}{\|\boldsymbol{r}-\boldsymbol{d}_{i}\|^{3}}+\dfrac{\boldsymbol{d}_{i}}{d_{i}^{3}}\right),$
$\boldsymbol{a}_{\mathrm{SRP}}(t,\boldsymbol{r})$ $=P_{\mathrm{Sun}}\left(\dfrac{\|\boldsymbol{d}_{\mathrm{Earth}}-\boldsymbol{d}_{\mathrm{Sun}}\|_{2}}{r_{\mathrm{Sun}}}\right)^{2}C_{r}\dfrac{A}{m}\dfrac{\boldsymbol{r}_{\mathrm{Sun}}}{r_{\mathrm{Sun}}},$
@@
The linearized solution to equation (1) in the vicinity of a reference state $\bar{\boldsymbol{\theta}}$ at time $t$ is given by
$\delta\boldsymbol{\theta}(t)=\dfrac{\partial\boldsymbol{\theta}(t)}{\partial\bar{\boldsymbol{\theta}}(t_{0})}\delta\boldsymbol{\theta}(t_{0}),$ (3)
where $\delta\boldsymbol{\theta}(t_{0})$ denotes the initial deviation from the initial reference state $\bar{\boldsymbol{\theta}}(t_{0})$ at time $t_{0}$. The partial ${\partial\boldsymbol{\theta}(t)}/{\partial\bar{\boldsymbol{\theta}}(t_{0})}$ is the state-transition matrix (STM) denoted by the shorthand $\boldsymbol{\Phi}(t,t_{0})$, obtained through the nonlinear integration of the initial value problem
$\begin{cases}\dot{\boldsymbol{\Phi}}\left(t,t_{0}\right)=\dfrac{\partial\boldsymbol{f}(t,\boldsymbol{\theta})}{\partial\bar{\boldsymbol{\theta}}}\boldsymbol{\Phi}\left(t,t_{0}\right),\\ \boldsymbol{\Phi}\left(t_{0},t_{0}\right)=\boldsymbol{I}_{6}.\end{cases}$ (4)
@@ 2.2 Near-Rectilinear Halo Orbit
Libration point orbits (LPO), periodic motions revolving around libration points, offer the option for spacecraft to be placed in unique periodic regimes in cislunar space that are otherwise unattainable by classical Keplerian orbits about the Earth or the Moon. There exist multiple families of LPOs, each with varying characteristics in terms of their spatial location, energy, and/or stability, and the NRHOs are a subset of the so-called halo orbit families about the L1/L2 points with bounded linear stability. For details on the dynamic structure of NRHOs, see [2, 3] and references therein. In this work, we focus on the 9:2 resonant NRHO of the Gateway, which completes 9 revolutions over 2 synodic periods of the Moon.
While periodic motions may be numerically constructed in simplified models such as the circular restricted three-body problem, the analogous motion in the full-ephemeris model is only quasi-periodic and contains unstable modes, thus requiring SK corrections to maintain the spacecraft on the desired orbit. Integrating motion along the NRHO is further complicated by the relatively low perilunes of around $3000\text{\,}\mathrm{k}\mathrm{m}$, as numerical integration errors get amplified during the transits through perilunes, where the dominant acceleration, proportional to $1/r^{2}$, has high sensitivities. The compounding effects of the inherent dynamical, as well as numerical, instabilities, necessitate a numerical integrator that is efficient with tight error tolerances. In this work, we make use of the explicit embedded Runge-Kutta Prince-Dormand 9(8) from Verner [23] in the GNU scientific library (GSL) [24].
@@ 2.3 System Architecture for Optical Navigation and Station-Keeping
Starting with the measurement block, synthetic images are generated using Blender; details of this implementation are provided in [21]. The image is then processed using the Christian-Robinson algorithm, which is a horizon-based, non-iterative method for generating a position measurement $\boldsymbol{r}_{\mathrm{P}}$, along with an analytical expression for the measurement covariance $\boldsymbol{P}_{\boldsymbol{r}_{\mathrm{P}}}$. The measurement and measurement covariance are fed to the navigation filter; for simulation purposes, we keep track of both the filter’s state estimate $\hat{\boldsymbol{\theta}}$ as well as the true state $\boldsymbol{\theta}$. The synthetic imaging block uses $\boldsymbol{\theta}$ to place the camera, while $\hat{\boldsymbol{\theta}}$ is used to point the camera toward the perceived location of the Moon. The control block on the bottom right is only given $\hat{\boldsymbol{\theta}}$ and $\boldsymbol{P}_{\boldsymbol{r}_{\mathrm{P}}}$ to compute a control command $\boldsymbol{u}$. The state estimation assumes that $\boldsymbol{u}$ is executed, while in fact a corrupted $\boldsymbol{u}+\delta\boldsymbol{u}$ is executed on the true state $\boldsymbol{\theta}$.

### v2

@@ Abstract
This article develops an optical navigation (OPNAV) and station-keeping pipeline for the near-rectilinear halo orbit (NRHO) in high-fidelity ephemeris model dynamics, using synthetic images in a non-iterative horizon-based OPNAV algorithm, applying the result in a navigation filter, and using the obtained estimates in a station-keeping control scheme that keeps the spacecraft in the vicinity of a reference orbit. We study differential correction-based and minimization-based implementations of the so-called $x$-axis crossing control scheme, and propose an improved targeting prediction scheme by incorporating the filter’s state covariance with an unscented transform. We also introduce a hysteresis mechanism, which improves station-keeping cost and provides insight into the difference in performance between the differential correction-based and minimization-based approaches. We perform Monte-Carlo experiments to assess the pipeline’s tracking and $\Delta V$ performances. We report several key findings, including the variability of the filter performance with the sensor field of view and measurement locations, station-keeping cost reduction achieved by the unscented transform-based prediction and hysteresis, as well as the variability of the cumulative $\Delta V$ as a function of maneuver location due to the periodic structure in the OPNAV-based filter’s estimation accuracy.
@@
$W^{(\ell)}_{m}$  = $\ell^{\rm th}$ sigma point weight for the mean
$W^{(\ell)}_{c}$  = $\ell^{\rm th}$ sigma point weight for the covariance
$\mathcal{X}^{(\ell)}$  = $\ell^{\rm th}$ sigma point
$\boldsymbol{x}$  = spacecraft inertial state vector
$\boldsymbol{y}$  = measurement vector
@@ 1 Introduction
A key challenge of operating space assets in cislunar space is the vast distance from the Earth. The distance makes frequent communication and navigational updates to cislunar spacecraft a challenge due to the scarcity and congestion of ground-based infrastructures, necessitating autonomous capability for guidance, navigation, and control (GN&C). A particular orbit of interest is the 9:2 resonant near-rectilinear halo orbit (NRHO) about the Earth-Moon L2 point [Lee2019, Zimovan-Spreen2020, Zimovan-Spreen2022], due to its planned use for the upcoming lunar space station, or Gateway, humanity’s next outpost in space.
@@
Regular station-keeping (SK) maneuvers, also referred to as orbit maintenance maneuvers, are required for spacecraft flying along the NRHO due to the orbit’s unstable modes. SK is required at an appropriate cadence in order to cancel the unstable modes in the presence of errors in state estimation, dynamics model, and control. As traffic on the NRHO is expected to increase, autonomous GN&C along this orbit is of critical importance, both to reduce reliance on ground-based navigation updates in nominal operations contexts, as well as to ensure a reliable onboard GN&C capability in case of off-nominal scenarios such as communication failures. Today, with the lack of a dedicated cislunar position, navigation, and timing (PNT) architecture, horizon-based optical navigation (OPNAV) [ChristianRobinson2016, Christian2017, Christian2021Tutorial] is a suitable measurement source that may be acquired and processed onboard a spacecraft along the NRHO. A number of prior simulation-based works considered autonomous navigation using horizon-based OPNAV on lunar libration point orbits (LPOs) [Franzese2019, Balossi2024], as well as specifically on the NRHO [Yun2020, Qi2023] and across various perspectives with respect to the Moon [Machuca2026-rc], but without consideration of SK activities. Several recent missions have successfully demonstrated OPNAV for orbit determination in cislunar flight: the EQUULEUS mission [Oguri2020-wi], the Orion spacecraft during the Artemis I mission [Inman2024], the LONEStar experiment during the Lunar Flashlight mission [Krause2024], and the CAPSTONE mission [Givens2025]. LUMIO [Cervone2022-ax] is another upcoming small spacecraft mission planned to use its optical payload for horizon-based OPNAV, among other purposes.
@@
There has been significant focus on SK algorithms and their performance under orbit determination error. Shirobokov et al. [Shirobokov2017] provides a survey of the various SK schemes and for LPOs in various dynamical regimes. Oguri [Oguri2024-of] demonstrates navigation with numerically simulated horizon-based OPNAV and SK based on an optimization-based pre-computed feedback policy. In the context of SK for NRHOs, the $x$-axis crossing control, where the $x$-axis corresponds to the Earth-Moon line, is known to perform favorably [Davis2017, Guzzetti2017, Newman2018, Davis2022]; in the recent CAPSTONE mission, this strategy was successfully deployed on the NRHO [Cheetham2021], and has been chosen as the SK scheme for the Gateway [Davis2022].
In this work, we evaluate the feasibility and performance of autonomous GN&C along NRHO by developing an end-to-end pipeline that consists of a navigation filter relying solely on measurements from processing realistic synthetic images of the Moon, and extensions to the $x$-axis crossing control algorithm that uses the recursively filtered state estimate to compute control actions. The contributions of this paper are as follows:
  • We consider the image processing for generating measurements via horizon-based OPNAV [ChristianRobinson2016, Christian2017] in parallel with navigation, where we augment the fidelity of the measurement covariance expression of the horizon-based OPNAV algorithm by accounting for the attitude error of the camera. We conduct Monte-Carlo experiments to report the levels of navigation accuracy that can be achieved by horizon-based OPNAV for various choices of camera, and quantify its limitations along the NRHO.
  • We compare two variants of the $x$-axis crossing control, namely the traditional differential correction-based approach [Guzzetti2017, Davis2017, Davis2022] and the optimization-based approach [Elango2022SequentialLinear]. In addition, we propose an unscented transform (UT)-based targeting scheme that uses the filter’s state covariance information to improve the state prediction. We also introduce a hysteresis mechanism within the $x$-axis crossing control scheme, where a tunable margin is introduced to desensitize the control scheme from noise by setting a tighter targeting tolerance than a triggering tolerance condition that prompts the control maneuver. We compare variants of the $x$-axis crossing control through Monte-Carlo experiments, and we highlight the significant performance difference that arises due to the short-sighted, single maneuver SK optimization as an approximation of the infinite-horizon cumulative SK cost. We also report the cost reduction achieved via UT-based targeting.
  • The full GN&C pipeline is used to evaluate the SK performance along NRHO navigating based on horizon-based OPNAV alone. We also study the sensitivity of the cumulative $\Delta V$ cost to the SK maneuver location, impacted by the periodic structure of the filter covariance achieved by horizon-based OPNAV.
@@ 2.1 Equations of Motion
$\dot{\boldsymbol{x}}=\begin{bmatrix}\dot{\boldsymbol{r}}\\ \dot{\boldsymbol{v}}\end{bmatrix}=\boldsymbol{f}(t,\boldsymbol{x})=\begin{bmatrix}\boldsymbol{v}\\ -\dfrac{\mu}{r^{3}}\boldsymbol{r}+\sum_{i}\boldsymbol{a}_{N_{i}}(t,\boldsymbol{r})+\boldsymbol{a}_{\mathrm{SRP}}(t,\boldsymbol{r})+\boldsymbol{a}_{\mathrm{SH}}(t,\boldsymbol{r})\end{bmatrix}$, (1)
where $\boldsymbol{r}\in\mathbb{R}^{3}$ is the position vector of the spacecraft with respect to an unforced particle at the origin of $\mathcal{F}_{\rm I}$, $\boldsymbol{v}\triangleq\dot{\boldsymbol{r}}\in\mathbb{R}^{3}$ is the rate of change of $\boldsymbol{r}$ in $\mathcal{F}_{\rm I}$, and $\boldsymbol{x}^{T}=[\boldsymbol{r}^{T},\boldsymbol{v}^{T}]\in\mathbb{R}^{6}$ is the state vector. In the acceleration terms on the right-hand side, $r=\|\boldsymbol{r}\|$ is the distance from the center of the Moon to the spacecraft, $\mu$ is the standard gravitational parameter of the Moon, $\boldsymbol{a}_{N_{i}}$ is the third-body perturbing acceleration of body $i$, $\boldsymbol{a}_{\mathrm{SRP}}$ is the perturbing acceleration due to SRP, and $\boldsymbol{a}_{\mathrm{SH}}$ is the perturbing acceleration due to the spherical harmonics terms of the Moon, In this work, we consider third-body perturbations due to the Earth and the Sun; in $\boldsymbol{a}_{\rm SH}$, we consider the term due to J2 coefficient. While the J2 term is included as a representative perturbation due to SH, the Moon has SH coefficients with similar orders of magnitude as $J_{2}$ at up to degree and order 3, which is to be included for higher-fidelity analyses. Their inclusion is straightforward, and we avoided it here for simplicity. The perturbation terms are given by
$\boldsymbol{a}_{N_{i}}(t,\boldsymbol{r})$ $=-\mu_{i}\left(\dfrac{\boldsymbol{r}-\boldsymbol{d}_{i}}{\|\boldsymbol{r}-\boldsymbol{d}_{i}\|^{3}}+\dfrac{\boldsymbol{d}_{i}}{d_{i}^{3}}\right),$ (2)
$\boldsymbol{a}_{\mathrm{SRP}}(t,\boldsymbol{r})$ $=P_{\mathrm{Sun}}\left(\dfrac{\|\boldsymbol{d}_{\mathrm{Earth}}-\boldsymbol{d}_{\mathrm{Sun}}\|_{2}}{r_{\mathrm{Sun}}}\right)^{2}C_{r}(A/m)\dfrac{\boldsymbol{r}_{\mathrm{Sun}}}{r_{\mathrm{Sun}}},$ (3)
@@
The linearized solution to equation (1) in the vicinity of a reference state $\bar{\boldsymbol{x}}$ at time $t$ is given by
$\delta\boldsymbol{x}(t)=\dfrac{\partial\boldsymbol{x}(t)}{\partial\bar{\boldsymbol{x}}(t_{0})}\delta\boldsymbol{x}(t_{0}),$ (5)
where $\delta\boldsymbol{x}(t_{0})$ denotes the initial deviation from the initial reference state $\bar{\boldsymbol{x}}(t_{0})$ at time $t_{0}$. The partial ${\partial\boldsymbol{x}(t)}/{\partial\bar{\boldsymbol{x}}(t_{0})}$ is the state-transition matrix (STM) denoted by the shorthand $\boldsymbol{\Phi}(t,t_{0})$, obtained through the nonlinear integration of the initial value problem
$\begin{cases}\dot{\boldsymbol{\Phi}}\left(t,t_{0}\right)=\dfrac{\partial\boldsymbol{f}(t,\boldsymbol{x})}{\partial\bar{\boldsymbol{x}}}\boldsymbol{\Phi}\left(t,t_{0}\right),\\ \boldsymbol{\Phi}\left(t_{0},t_{0}\right)=\boldsymbol{I}_{6}.\end{cases}$ (6)
@@ 2.2 Near-Rectilinear Halo Orbit
Libration point orbits (LPO), periodic motions revolving around libration points, offer the option for spacecraft to be placed in unique periodic regimes in cislunar space that are otherwise unattainable by classical Keplerian orbits about the Earth or the Moon. There exist multiple families of LPOs, each with varying characteristics in terms of their spatial location, energy, and/or stability, and the NRHOs are a subset of the so-called halo orbit families about the L1/L2 points with bounded linear stability. For details on the dynamic structure of NRHOs, see [Zimovan-Spreen2020, Zimovan-Spreen2022] and references therein. In this work, we focus on the 9:2 resonant NRHO of the Gateway, which completes 9 revolutions over 2 synodic periods of the Moon.
While periodic motions may be numerically constructed in simplified models such as the circular restricted three-body problem, the analogous motion in the full-ephemeris model is only quasi-periodic and contains unstable modes, thus requiring SK corrections to maintain the spacecraft on the desired orbit. Integrating motion along the NRHO is further complicated by the relatively low perilunes of around $3000\text{\,}\mathrm{k}\mathrm{m}$, as numerical integration errors get amplified during the transits through perilunes, where the dominant acceleration, proportional to $1/r^{2}$, has high sensitivities. The compounding effects of the inherent dynamical, as well as numerical, instabilities, necessitate a numerical integrator that is efficient with tight error tolerances. In this work, we make use of the explicit embedded Runge-Kutta Prince-Dormand 9(8) from Verner [Verner2010] in the GNU scientific library (GSL) [gough2009gnu].
@@ 2.3 System Architecture for Optical Navigation and Station-Keeping
Starting with the measurement block, synthetic images are generated using Blender; details of this implementation are provided in [Shimane2023oct]. The image is then processed using the Christian-Robinson algorithm, which is a horizon-based, non-iterative method for generating a position measurement $\boldsymbol{r}_{\mathrm{P}}$, along with an analytical expression for the measurement covariance $\boldsymbol{P}_{\boldsymbol{r}_{\mathrm{P}}}$. The measurement and measurement covariance are fed to the navigation filter; for simulation purposes, we keep track of both the filter’s state estimate $\hat{\boldsymbol{x}}$ as well as the true state $\boldsymbol{x}$. The synthetic imaging block uses $\boldsymbol{x}$ to place the camera, while $\hat{\boldsymbol{x}}$ is used to point the camera toward the perceived location of the Moon. The control block on the bottom right is only given $\hat{\boldsymbol{x}}$ and $\boldsymbol{P}_{\boldsymbol{r}_{\mathrm{P}}}$ to compute a control command $\boldsymbol{u}$. The state estimation assumes that $\boldsymbol{u}$ is executed, while in fact a corrupted $\boldsymbol{u}+\delta\boldsymbol{u}$ is executed on the true state $\boldsymbol{x}$.